SOLVE:
3x^2 - 41x = -60

Answer by mananth(16946)

3x^2 - 41x = -60
..
3x^2 - 41x+60=0
..
3x^2-36x-5x+60=0
3x(x-12)-5(x-12)=0
(x-12)(3x-5)=0
x =12 OR x= 5/3
...
m.ananth@hotmail.ca